OpenStudy (anonymous):

The deposits Ginny makes at her bank each month form an arithmetic sequence. The deposit for month 3 is $150, and the deposit for month 5 is $180. Answer the questions below and show all work. 1. What is the common difference for the deposits made each month? Somebody please show me how I can solve this? I'm pretty sure it is 15, but I need to show work somehow.

OpenStudy (anonymous):

@JamesJ @satellite73 @campbell_st @amistre64

OpenStudy (campbell_st):

ok... so you need to find the common difference in a 2 month period the value of the deposit increases by $30 so $30/2 is the monthly increase in the deposit or the common difference.

OpenStudy (campbell_st):

if you need an algebraic solution then a term in an arithmetic sequence is \[T_{n} = a + (n - 1) d\] a = 1st term, n = number of terms and d = common difference so \[150 = a + (3 -1)d \] \[180 = a + ( 5 - 1)d\] simplifying you have Equation 1 \[150 = a + 2d\] equation 2 \[180 = a + 4d \] solve simultaneously by elimination Equation2 - Equation 2 a + 4d = 180 - a + 2d = 150 ---------------- 2d = 30 d = 15

OpenStudy (anonymous):

Thanks, the formula is just what I was looking for! Could you possibly help me write and explicit formula for the sequence now? @campbell_st

OpenStudy (campbell_st):

so if you know d then you need to find a the 1st deposit or 1st term \[150 = a + (3 -1)15\] a = 120 so the formula for any deposit is \[T_{n} = 120 + (n -1)15\] just simplify the formula above

OpenStudy (anonymous):

So \[a _{n}=15n+105\] ?

OpenStudy (campbell_st):

thats about it... well done..

OpenStudy (anonymous):

Thanks a ton!